How many different 5-card poker hands can be made from a deck of 52 cards?
Two equal circles are cut out of a rectangle of dimensions 16 by 8. The circles have the maximum diameter possible. What is the approximate area of the paper remaining after the circles have been cut out?

Answer by stanbon(75887)

How many different 5-card poker hands can be made from a deck of 52 cards?
----
52C5 = 52!/[(52-5)!*5!] = [52*51*50*49*48]/[1*2*3*4*5] = 2598960
----------------------------------------

Two equal circles are cut out of a rectangle of dimensions 16 by 8. The circles have the maximum diameter possible. What is the approximate area of the paper remaining after the circles have been cut out?
----
Area of the rectangle: 16*8 = 128 sq units
----
Area of each circle: (pi)*4^2= 16(pi)
----
Area of remaining paper: 128-2*16(pi) = 27.47 sq units